Handout for Assignment 7.1: Bond Valuation

Use the following data to answer the assignment questions.

Assume you are evaluating whether to purchase the following $1,000 face value bonds:

•   Co. X bond with a 6% coupon rate that matures in 9 years.

•   Co. Y bond with an 11% coupon rate that matures in 7 years.

Answer the following questions:
Use a spreadsheet file to calculate and report the following information:

  • Value these bonds assuming a market rate on similar risk bonds is 7% and interest is paid annually.
  • Value these bonds assuming a market rate on similar risk bonds is 7% and interest is paid semi-annually.
  • Value these bonds assuming a market rate on similar risk bonds is 12% and interest is paid annually.
  • Assuming both bonds were issued at the same time, why would the Co. Y bond pay a higher coupon rate?
Answer #1

FACE Value coupon payment Period (in years) 1000 1000 110 L 97 60 9 10 11 X Y 12 7% 7% Particular Interest rate (in market) N

(a) Value of Bond =-PV(C12,C7,C6,C5)

(b) Value of Bond =-PV(C16/C17,C7*C17,C6/C17)-PV(C16,C7,0,C5)

(c) Value of Bond =-PV(C20,C7,C6,C5)

(d) Value Of Bond=-PV(C24/C25,C7*C25,C6/C25)-PV(C24,C7,0,C5)

​​​​​​Y Co. Pays higher rate because it might have Higher risk of default as compared to X LTD. It is called Default risk Premium.

I hope my efforts will be fruitful to you...?